Assume that you are a consultant to Broske Inc., and you have been provided with the following data: D1 = $0.67; P0 = $47.50; and g = 8.00% (constant). What is the cost of equity from retained earnings?
Answer #1

Given,

D1 = $0.67

P0 = $47.50

g = 8.00% or 0.08


Cost of equity from retained earnings = D1/P0 + g

= $0.67/$47.50 + 0.08

= 0.0141 + 0.08 = 0.0941 


answered by: gavin